40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat May 14 08:36:03 2022, Chris, Update, DAQ, DAQ troubleshooting 
    Reply  Mon May 16 10:49:01 2022, Anchal, Update, DAQ, DAQ troubleshooting 
       Reply  Mon May 16 12:59:27 2022, Chris, Update, DAQ, DAQ troubleshooting timeseries.pngerr.png
Message ID: 16854     Entry time: Mon May 16 10:49:01 2022     In reply to: 16853     Reply to this: 16855
Author: Anchal 
Type: Update 
Category: DAQ 
Subject: DAQ troubleshooting 

[Anchal, Paco, JC]

Thanks Chris for the fix. We are able to access the testpoints now but we started facing another issue this morning, not sure how it is related to what you did.

  • The C1:LSC-TRX_OUT and C1:LSC-TRY_OUT channels are stuck to zero value.
  • These were the channels we used until last friday to align the interferometer.
  • These channels are routed through the c1rfm FE model (Reflected Memory model is the name, I think). These channels carry the IR transmission photodiode monitors at the two ends of the interferometer, where they are first logged into the local FEs as C1:SUS-ETMX_TRX and C1:SUS-ETMY_TRY .
  • These channels are then fed to C1:SCX-RFM_TRX -> C1:RFM_TRX -> C1:RFM-LSC_TRX -> C1:LSC-TRX and similar for Y side.
  • We are able to see channels in the end FE filtermodule testpoints (C1:SUS-ETMX_TRX_OUT & C1:SUS-ETMY_TRY_OUT)
  • However, we are unable to see the same signal in c1rfm filter module testpoints like C1:RFM_TRX_IN1, C1:RFM_TRY_IN1 etc
  • There is an IPC error shown in CDS FE status screen for c1rfm in c1sus. But we remember seeing this red for a long time and have been ignoring it so far as everything was working regardless.

The steps we have tried to fix this are:

  • Restart all the FE models in c1lsc, c1sus, and c1ioo (without restarting the computers themselves) , and then burt restore.
  • Restart all the FE models in c1iscex, and c1iscey (only c1iscey computer was restarted) , and then burt restore.

These above steps did not fix the issue. Since we have  the testpoints (C1:SUS-ETMX_TRX_OUT & C1:SUS-ETMY_TRY_OUT) for now to monitor the transmission levels, we are going ahead with our upgrade work without resovling this issue. Please let us know if you have any insights.

ELOG V3.1.3-